C=50+0.80 (Y-T)
I=200
G=100
Find out the equilibrium level of income
Suppose G increase to 125 what is the new equilibrium level of income?
What level of G is needed to achieve a target income of 200?
1
Expert's answer
2020-05-24T18:09:41-0400

Ye = C + I + G, 

The equilibrium level of income becomes; Ye=50+0.80 (Y-T) +200+ 100

Ye=350+ 0.80(Y-T)

If G increases to 125 , then Ye=50+0.80 (Y-T) +200+ 125

Ye=375+ 0.80(Y-T)

If the income is 200, then G becomes

200=50+0.80 (Y-T) +200+ G

G=-50-0.80(Y+T)